OpenStudy (anonymous):

The steps in the simplification of the expression 46 + 7x + 4y −2x – y − 27 are shown below. Step 1: 46 – 27 + 7x − 2x + 4y − y Step 2: (46 − 27) + (7x − 2x) + (4y − y) Step 3: 19 + x (7 − 2) + y (4 − 1) Step 4: 19 + x (5) + y (3) Step 5: 19 + 5x + 3y Which statement is correct?

OpenStudy (anonymous):

Please help. :l I'm lost.

OpenStudy (anonymous):

Step 1 uses commutative property of addition, and step 5 uses commutative property of multiplication. Step 4 uses associative property of multiplication, and step 3 uses simplification. Step 2 uses associative property of addition, and step 5 uses simplification. Step 1 uses distributive property of addition, and step 3 uses distributive property of multiplication

jimthompson5910 (jim_thompson5910):

are you familiar with terms like commutative property of addition or associative property of addition?

OpenStudy (anonymous):

yes. But I can't figure this out. I even looked it up.

jimthompson5910 (jim_thompson5910):

commutative property of addition: you can add in any order, so x+y = y+x or 1+2 = 2+1 associative property of multiplication: you can rearrange the parenthesis when multiplying example: 2*(3*4) = (2*3)*4 associative property of addition: you can rearrange the parenthesis when adding example: 2+(3+4) = (2+3)+4

jimthompson5910 (jim_thompson5910):

going from 46 + 7x + 4y −2x – y − 27 to 46 – 27 + 7x − 2x + 4y − y what happened?

OpenStudy (anonymous):

It would be commutative then?

jimthompson5910 (jim_thompson5910):

yes because we're swapping the terms (since you can add those terms in any order)